OpenStudy (anonymous):

Solve the following equation. 3x² - 10x + 8 =0 Choose one answer. a. x = 4/3, 1 b. x= 3,10 c. x = 3/2, 4 d. no solution

OpenStudy (anonymous):

3x^2-6x-4x+8=0 or 3x(x-2)-4(x-2)=0 or (3x-4)(x-2)=0 either 3x-4 = 0 or x-2 = 0 ie either x=4/3 or x= 2

OpenStudy (anonymous):

So, none of these options are correct. Therefore the answer is d?

OpenStudy (anonymous):

to say that the equation has no solution would be incorrect. as clearly x=4/3 and x =2 are the solutions. instead of giving no solutions as an option, the problem should have given none of the above as option d

OpenStudy (anonymous):

Ahhh. Yes, agreed. Thank you :)

OpenStudy (anonymous):

ur welcome :)